OpenStudy (anonymous):

Jacqueline drove 72 km on 6 L of gasoline. Which of the following equations can be used to find how far she can travel on a full tank of 50 L?

OpenStudy (anonymous):

600 miles but how would you write that..

OpenStudy (anonymous):

If Jacqueline drove 72 km on 6 L of gasoline, she should be able to travel 600 miles on a full 50 L tank. Divide 50 L (full tank) by 6 L (known variable). Multiply 8.33 * 72 (known variable) to get the final answer of 600 miles.

OpenStudy (anonymous):

you have a ratio of 72/6. if you simplify that you will get a ratio of 12/1. now you only have to multiply 12 times 50 to find the answer. that is in fact 600 \[\frac{ 72 }{ 6 } \frac{ x }{ 50 }\] after the simplification you get \[\frac{ 12 }{ 1}\frac{ x }{ 50}\] now you cross multiply to get the answer.